OpenStudy (anonymous):

Find the solution to this exponential equations, correct to four decimal places. 10^(1-x) = 6^x

OpenStudy (anonymous):

help pleasee

OpenStudy (anonymous):

\[10^{(1-x)}=6^{x}\]\[\log_{10}10^{(1-x)}=\log_{10}6^{x} \]\[1-x=x \log_{10}6 \]\[(1-x)/x=\log_{10}6 \]\[1/x-1=\log_{10}6 \]\[1/x=(\log_{10}6)+1 \]\[x=1/[(\log_{10}6)+1]\]\[x \approx.562\]